# Predicting Grain Boundary Segregation in Magnesium Alloys: An Atomistically Informed Machine Learning Approach

### v1

@@ III.1 Thermal properties of pure Mg
To benchmark the performance of the MEAM potential in predicting thermal properties of pure Mg, the temperature-dependent vibrational heat capacities at constant volume $C_{\text{v}}$ and pressure $C_{\text{p}}$ were calculated according to Equations 10 and 11 and compared with ab-initio [54] and experimental [55] data, see Figure 1(a). The calculated temperature-dependent $C_{\text{v}}$ using the MEAM potential shows excellent agreement with the ab-initio result. In addition, the calculated $C_{\text{p}}$ using the MEAM potential aligns closely with the ab-initio data and experimental measurements. Overall, these comparisons demonstrate that the MEAM potential effectively describes the thermal properties of pure Mg.
The per-site vibrational free energies $F^{\text{vib}}$ for over 38,000 GB sites within the Mg polycrystal, as characterized using the adaptive CNA method, were calculated following Equation 8. The distribution of $F^{\text{vib}}$ at 600 K across these GB sites is shown in Figure 1(b), fitted with a skew-normal function characterized by the following parameters: location $\xi$= --0.177 eV/atom, scale $\omega$= 0.001 eV/atom, and shape $\alpha$= 0.047, leading to a mean value $\mu$ of --0.176 eV/atom. The small $\alpha$ value alongside the closely matched $\xi$ and $\mu$ values suggests that the distribution is nearly symmetric, closely approximating a normal distribution. Notably, a clear deviation from the $F^{\text{vib}}$ value of the bulk site at 600 K (--0.1743 eV/atom) was identified, highlighting significant differences in thermal properties between GB and bulk sites.
@@ III.2 Segregation free energy at finite temperatures
Table 1: Summary of fitting parameters, including location ($\xi$, in eV/atom), scale ($\omega$, in eV/atom), and shape ($\alpha$, dimensionless) parameters, as well as mean value ($\mu=\xi+\omega\frac{\alpha}{\sqrt{1+\alpha^{2}}}\sqrt{\frac{2}{\pi}}$ in eV/atom) for skew-normal distribution applied to segregation energy $\Delta E_{\text{seg}}$ at 0 K, segregation vibrational free energy $\Delta F_{\text{seg}}^{\text{vib}}$ and segregation free energy $\Delta F_{\text{seg}}$ at 600 K in Mg-M (M=Nd, Ca, Y, Li, Al and Zn) alloy systems.
@@
The per-site segregation energies $\Delta E_{\text{seg}}$ for solute elements (Nd, Ca, Y, Li, Al, and Zn) at 0 K of GB sites within the Mg polycrystal were calculated. The elements Nd, Ca, and Y are considered larger solutes relative to the Mg atom, as substituting Mg with these elements induces positive volumetric strain in the Mg lattice. In contrast, the elements Li, Al, and Zn, having smaller atomic radii, introduce negative volumetric strain to the Mg matrix upon substitution [33]. The resulting $\Delta E_{\text{seg}}$ distributions in these systems are illustrated in Figure 2. For all simulated Mg alloys in this work, the $\Delta E_{\text{seg}}$ spectra exhibit a skew-normal distribution, similar to observations in FCC polycrystalline systems as reported by [65, 35]. A skew-normal function was applied to fit the $\Delta E_{\text{seg}}$ spectra, with the fitting parameters $\xi$, $\omega$ and $\alpha$ detailed in TABLE 1. The majority of the GB sites across all studied Mg alloy systems are energetically favorable to solute segregation, as evidenced by their negative $\Delta E_{\text{seg}}$. This trend is further underscored by the mean ($\mu$) values of the $\Delta E_{\text{seg}}$ spectra, which are also negative for all systems. Notably, Nd shows the most negative $\mu$ value of --0.154 eV/atom, with Li and Al following with $\mu$ of --0.100 and --0.081 eV/atom, respectively. For the Nd, Ca, Y, and Li solutes, the distribution of $\Delta E_{\text{seg}}$ exhibits a left-skewed pattern with the mean being pulled toward the left tail, reflected by negative $\alpha$ values. In contrast, the Al and Zn solutes demonstrate a right-skew distribution, as indicated by positive $\alpha$ values, suggesting that the mean of these distributions is pulled towards the right tail. Among the solute elements studied, Nd exhibits the broadest distribution of $\Delta E_{\text{seg}}$ followed by the Ca element, as indicated by the values of $\omega$.
@@
Figure 3: Grain boundary concentration of M solutes (M=Nd, Ca, Y, Li, Al, and Zn) as a function of temperature with and without considering the vibrational entropy contribution at (a) 0.1 at.and (b) 1 at.bulk solute concentrations following the Langmuir–McLean model, as denoted in Equations 12 and 13. The shaded areas indicate the differences between predicted GB solute concentrations with and without considering the vibrational entropy contribution.
@@
The segregation vibrational free energies ($\Delta F_{\text{seg}}^{\text{vib}}$) at finite temperatures were calculated by the HA approach according to Equation 5. The distribution of $\Delta F_{\text{seg}}^{\text{vib}}$ follow a skew-normal function in all simulated Mg alloy systems, similar to the $\Delta E_{\text{seg}}$ distributions. The spread of the $\Delta F_{\text{seg}}^{\text{vib}}$ distributions at 600 K is significantly narrower than that of $\Delta E_{\text{seg}}$, with $\omega$ values spanning from 0.044 to 0.074 eV/atom. Notably, Y not only presents the narrowest distribution within the $\Delta E_{\text{seg}}$ spectrum but also exhibits the most constrained spread in the $\Delta F_{\text{seg}}^{\text{vib}}$ spectrum. The $\mu$ values are slightly positive for all alloy systems except for Al, which exhibits a $\mu$ value of --0.001 eV/atom. In addition, Al is distinguished as the only element exhibiting a negative $\alpha$ value of --0.801 in the $\Delta F_{\text{seg}}^{\text{vib}}$ spectra, demonstrating a left-skewed distribution where the mean is pulled towards the left tail.
@@
The segregation free energies $\Delta F_{\text{seg}}$ at finite temperatures were determined by adding the vibrational entropy contribution to the segregation energy $\Delta E_{\text{seg}}$ (Equation 4). The $\Delta F_{\text{seg}}$ distributions of the Mg alloy systems were fitted using a skew-normal function as shown in Figure 2. Compared to the $\Delta E_{\text{seg}}$ distributions, a shift towards positive values was noted for all solute elements, implying a reduced tendency for segregation with the increasing temperature. Despite this shift, all $\mu$ values of the $\Delta F_{\text{seg}}$ distribution remained negative at 600 K, with Nd showing the most negative value at --0.127 eV/atom. For the Al element, the transition in segregation behavior with temperature is less marked (see Figure 2(e)), as the $\mu$ value of the $\Delta F_{\text{seg}}$ distribution remains consistent with that of the $\Delta E_{\text{seg}}$ distribution, both measured at --0.081 eV/atom. A contraction in the distributions of $\Delta F_{\text{seg}}$ was evident from the decrease in the $\omega$ value. Notably, the distribution for Ca demonstrated a significant change in skewness at 600 K, with the skewness parameter $\alpha$ shifting from --0.669 at 0 K to a positive value of 0.488.
@@
Besides quantifying energy spectra, the per-site segregation (free) energy, denoted as $\Delta G_{i}$, can be directly correlated with solute concentration at GBs based on the Langmuir–McLean model, as illustrated in Equation 12 in the context of an infinitesimally dilute solid solution:
@@
where $X_{\text{bulk}}$ is the solute concentration in the bulk region, and $X_{\text{GB},i}$ is the fraction of solute concentration of a GB site $i$. This correlation facilitates the calculation of the weighted average solute concentration at GBs ($X_{\text{GB}}$) as described in Equation 13:
@@
The effect of vibrational entropy on solute segregation at GBs was illustrated by calculating the fractional solute concentration of individual sites according to Equation 12 with and without considering the vibrational term $\Delta F_{\text{seg}}^{\text{vib}}$. The resulting concentrations of various solutes (Nd, Ca, Y, Li, Al, and Zn) at GBs within the simulated polycrystalline Mg at $X_{\text{bulk}}$ of 0.1 at.and 1 at.as a function of temperature are shown in Figure 3. Accounting for the vibrational entropy contribution results in notably lower solute concentrations at GBs than those derived from calculations relying solely on $\Delta E_{\text{seg}}$. The difference in solute concentrations at GBs, attributed to vibrational entropy, becomes increasingly significant with rising temperature and $X_{\text{bulk}}$. Among the solutes studied, Nd stands out by exhibiting a significantly higher $X_{\text{GB}}$ value across the temperature range up to 800 K examined. This range encompasses the conventional thermomechanical processing temperatures of Mg alloys in experimental settings.
@@ III.3 Machine learning prediction of segregation energy and segregation free energy
Figure 4: Contour plots of ML predicted versus true (a-c) per-site segregation energy $\Delta E_{\text{seg}}$ and (d-f) segregation free energy $\Delta F_{\text{seg}}$ at 600 K of Zn solute segregation at GB sites in the Mg polycrystal. Only the validation data are illustrated here. The CatBoostRegressor model was trained separately using the (a,c) ESD features, (b,e) SOAP features, and (c,f) both ESD and SOAP features. The color density indicates the concentration of data points, with denser regions shown in lighter shades. The diagonal red dashed line represents the ideal scenario where predicted values match the true values perfectly.
@@
Estimating the solute segregation tendency, particularly calculating the vibrational entropy contribution via the HA approach, is computationally expensive for all GB sites in polycrystals. The ML approach is ideal for developing a model to predict the segregation tendency of solutes to GB sites and capture the segregation energy spectrum based on accessible energetic and structural features, which are computationally inexpensive. To predict $\Delta E_{\text{seg}}$, $\Delta F_{\text{seg}}^{\text{vib}}$, and $\Delta F_{\text{seg}}$ at 600 K, we generated 16 ESD that capture the local atomic environment and thermal vibrational properties of GB sites in the polycrystal. The Pearson’s correlation matrices of these features with target segregation energies in different Mg alloying systems are illustrated in Figures S4-9 to identify which features have strong relationships with the target variables and with each other. Among all ESD, $\sigma_{h}$, which is calculated based on $\sigma_{xx}$, $\sigma_{yy}$, and $\sigma_{zz}$, shows the highest correlation coefficient above 0.80 with $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ at 600 K for all solute elements, indicating its significant role in determining the segregation behavior of solute atoms in the material. The normal stress components $\sigma_{xx}$, $\sigma_{yy}$, and $\sigma_{zz}$ exhibit similar correlation coefficients. Additionally, potential energy, atomic volume, surface area, and coordination number also exhibit high correlation coefficients with segregation energies.
@@
The contour plots in Figure 4(a) reveal that the predicted $\Delta E_{\text{seg}}$ values align closely with the true values along the diagonal red dashed line, which represents the ideal prediction scenario. The color density indicates the concentration of data points, with denser regions shown in brighter shades, highlighting that the highest concentration of data points is near the ideal prediction line. This indicates the CatBoostRegressor model has a strong predictive capability with the ESD feature set. The good performance of the CatBoostRegressor model in predicting $\Delta E_{\text{seg}}$ with the ESD features is further demonstrated by the $R^{2}$ scores, with a minimal $R^{2}$ of 0.82 for the Nd solute element and above 0.90 for Ca, Li, and Zn solute elements, see Figure 4 and Figure S10. In addition, the distributions of the ESD exhibit a high degree of overlap between the predicted and the calculated $\Delta E_{\text{seg}}$, as shown in Figure 5 and Figure S11 for all solute elements. Among all ESD features, $\sigma_{h}$ demonstrates the highest importance, except for the Li solute element, where potential energy was identified as the most important feature, see Figure S12. Potential energy was ranked as the second most important feature for Al and Zn solute elements, and third for Ca. The structural features obtained using Voronoi tessellation, including atomic volume, surface area, and cavity radius, were ranked highly in importance, e.g., for Nd and Zn solute elements, atomic volume is the second most important feature, while for the Y solute element, cavity radius was ranked second. The vibrational properties calculated from the series of MD simulations, including $MSD_{\text{vib}}$ and $V_{\text{flex}}$, were demonstrated as important features in predicting $\Delta E_{\text{seg}}$: for Nd, Ca, Y, Al, and Zn, these vibrational features are among the top five most important features.
@@
Figure 5: Distributions of ML predicted and true (a-c) per-site segregation energy $\Delta E_{\text{seg}}$ and (d-f) segregation free energy $\Delta F_{\text{seg}}$ at 600 K of Zn solute segregation at GB sites in the Mg polycrystal. Only the validation data are illustrated here. The CatBoostRegressor model was trained separately using the (a,c) ESD features, (b,e) SOAP features, and (c,f) both ESD and SOAP features. Distributions of predicted versus true energies with a bin size of 0.01 eV/atom.
@@
In addition to the ESD features, we generated SOAP features for automatically describing the fingerprint of the local atomic environment at GB sites in the polycrystal. These SOAP descriptors were fed to the CatBoostRegressor model to predict $\Delta E_{\text{seg}}$, $\Delta F_{\text{seg}}^{\text{vib}}$, and $\Delta F_{\text{seg}}$ at 600 K (Figure 4-5 and Figure S19-24). The SOAP-featured model shows better performance in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ at 600 K than the ESD-featured model for all solute elements, except for Li, as demonstrated by higher $R^{2}$ scores and lower MAE (see Figure 6). However, the performance in predicting $\Delta F_{\text{seg}}^{\text{vib}}$ remains poor (Figure S21,22). By merging ESD and SOAP descriptors, the ML model shows further improvement in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$, as evidenced by the tighter clustering around the diagonal (Figure 4(c,f)) and with the $R^{2}$ score of all solute element in predicting $\Delta E_{\text{seg}}$ above 0.85 and $\Delta F_{\text{seg}}$ above 0.80 (Figure S25-30). However, the improvement by utilizing combined ESD and SOAP features in predicting $\Delta F_{\text{seg}}^{\text{vib}}$ is limited.
@@
The residual distributions of predicted and true $\Delta E_{\text{seg}}$, $\Delta F_{\text{seg}}^{\text{vib}}$, and $\Delta F_{\text{seg}}$ via ESD, SOAP, and combined ESD and SOAP features are shown in Figure S31-33. A left-skewed residual distribution in $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ was observed for all solute elements and feature sets, indicating the models consistently over or under-predict the solute segregation tendency at favorable segregation sites. For future work, uncertainty determination will be crucial to better understand and address these biases and refine the models accordingly.

### v2

@@ III.1 Thermal properties of pure Mg
To benchmark the performance of the MEAM potential in predicting thermal properties of pure Mg, the temperature-dependent vibrational heat capacities at constant volume $C_{\text{v}}$ and pressure $C_{\text{p}}$ were calculated (see Section II) and compared with ab-initio [68] and experimental [69] data, see Figure 1(a). The calculated temperature-dependent $C_{\text{v}}$ using the MEAM potential shows excellent agreement with the ab-initio result. In addition, the calculated $C_{\text{p}}$ using the MEAM potential aligns closely with the ab-initio data and experimental measurements. Overall, these comparisons demonstrate that the MEAM potential effectively describes the thermal properties of pure Mg.
Figure 1: (a) The vibrational heat capacities at constant volume $C_{\text{v}}$ and pressure $C_{\text{p}}$ of bulk Mg (containing 5 $\times$ 5 $\times$ 5 orthogonal unit cells) as functions of temperature, calculated by the full harmonic approximation (HA) method using the MEAM potential. For comparison, results from density functional perturbation theory (DFPT) [68] and experimental data [69] are presented. (b) The distribution of $F^{\text{vib}}$ of all GB sites in the Mg polycrystal at 600 K calculated based on HA ($r_{\text{HA}}$=14 Å, bin size: 1×$\times 10^{-4}$ eV/atom). A skew-normal function (dashed curve) was fitted to the $F^{\text{vib}}$ distribution. The dashed line indicates the $F^{\text{vib}}$ value for bulk Mg.
The per-site vibrational free energies $F^{\text{vib}}$ for over 38,000 GB sites within the Mg polycrystal, as characterized using the adaptive CNA method, were calculated following Section II. The distribution of $F^{\text{vib}}$ at 600 K across these GB sites is shown in Figure 1(b), fitted with a skew-normal function characterized by the following parameters: location $\xi$= --0.177 eV/atom, scale $\omega$= 0.001 eV/atom, and shape $\alpha$= 0.047, leading to a mean value $\mu$ of --0.176 eV/atom. The small $\alpha$ value alongside the closely matched $\xi$ and $\mu$ values suggests that the distribution is nearly symmetric, closely approximating a normal distribution. Notably, a clear deviation from the $F^{\text{vib}}$ value of the bulk site at 600 K (--0.1743 eV/atom) was identified, highlighting significant differences in thermal properties between GB and bulk sites.
@@ III.2 Segregation free energy at finite temperatures
Table 1: Summary of fitting parameters, including location ($\xi$, in eV/atom), scale ($\omega$, in eV/atom), and shape ($\alpha$, dimensionless) parameters, as well as mean value ($\mu=\xi+\omega\frac{\alpha}{\sqrt{1+\alpha^{2}}}\sqrt{\frac{2}{\pi}}$, in eV/atom) for skew-normal distribution applied to segregation energy $\Delta E_{\text{seg}}$ at 0 K, segregation vibrational free energy $\Delta F_{\text{seg}}^{\text{vib}}$ and segregation free energy $\Delta F_{\text{seg}}$ at 600 K in Mg-X (X=Nd, Ca, Y, Li, Al and Zn) alloy systems.
@@
The per-site segregation energies $\Delta E_{\text{seg}}$ for solute elements (Nd, Ca, Y, Li, Al, and Zn) at 0 K of GB sites within the Mg polycrystal were calculated. The elements Nd, Ca, and Y are considered larger solutes relative to the Mg atom, as substituting Mg with these elements induces positive volumetric strain in the Mg lattice. In contrast, the elements Li, Al, and Zn, having smaller atomic radii, introduce negative volumetric strain to the Mg matrix upon substitution [38]. The resulting $\Delta E_{\text{seg}}$ distributions in these systems are illustrated in Figure 2. For all simulated Mg alloys in this work, the $\Delta E_{\text{seg}}$ spectra exhibit a skew-normal distribution, similar to observations in FCC polycrystalline systems as reported by [70, 40]. A skew-normal function was applied to fit the $\Delta E_{\text{seg}}$ spectra, with the fitting parameters $\xi$, $\omega$ and $\alpha$ detailed in Table 1. The majority of the GB sites across all studied Mg alloy systems are energetically favorable to solute segregation, as evidenced by their negative $\Delta E_{\text{seg}}$. This trend is further underscored by the mean ($\mu$) values of the $\Delta E_{\text{seg}}$ spectra, which are also negative for all systems. Notably, Nd shows the most negative $\mu$ value of --0.154 eV/atom, with Li and Al following with $\mu$ of --0.100 and --0.081 eV/atom, respectively, highlighting their strong segregation tendencies.
The skewness parameter ($\alpha$) reflects the asymmetry in segregation energy distributions. For the Nd, Ca, Y, and Li solutes, the distribution of $\Delta E_{\text{seg}}$ exhibits a left-skewed pattern with the negative mean being pulled toward the left tail, as evidenced by negative $\alpha$ values. This indicates a strongly favorable segregation tendency, as both the central tendency and extreme tails favor segregation. On the other hand, the Al and Zn solutes demonstrate a right-skew distribution, characterized by positive $\alpha$ values, indicating that the negative mean of these distributions is shifted towards the right tail. This suggests mixed segregation behavior, where most sites are favorable, but a significant number of less favorable sites exist. Among the solute elements studied, Nd exhibits the broadest distribution of $\Delta E_{\text{seg}}$ followed by Ca, as indicated by the values of $\omega$, which suggests greater variability in segregation behavior across GB sites.
@@
Figure 3: Grain boundary concentration of X solutes (X=Nd, Ca, Y, Li, Al, and Zn) as a function of temperature with and without considering the vibrational entropy contribution at (a) 0.1 at.and (b) 1 at.bulk solute concentrations following the Langmuir–McLean model, as denoted in Equations (12) and (13). The shaded areas indicate the differences between predicted GB solute concentrations with and without considering the vibrational entropy contribution.
@@
The segregation vibrational free energies $\Delta F_{\text{seg}}^{\text{vib}}$ at finite temperatures were calculated by the HA approach according to Section II. The distribution of $\Delta F_{\text{seg}}^{\text{vib}}$ follow a skew-normal function in all simulated Mg alloy systems, similar to the $\Delta E_{\text{seg}}$ distributions. The spread of the $\Delta F_{\text{seg}}^{\text{vib}}$ distributions at 600 K is significantly narrower than that of $\Delta E_{\text{seg}}$, with $\omega$ values spanning from 0.044 to 0.074 eV/atom. Notably, Y not only presents the narrowest distribution within the $\Delta E_{\text{seg}}$ spectrum but also exhibits the most constrained spread in the $\Delta F_{\text{seg}}^{\text{vib}}$ spectrum. The $\mu$ values are slightly positive for all alloy systems except for Al, which exhibits a $\mu$ value of --0.001 eV/atom. In addition, Al is distinguished as the only element exhibiting a negative $\alpha$ value of --0.801 in the $\Delta F_{\text{seg}}^{\text{vib}}$ spectra, demonstrating a left-skewed distribution where the mean is pulled towards the left tail.
@@
The segregation free energies $\Delta F_{\text{seg}}$ at finite temperatures were determined by adding the vibrational entropy contribution to the segregation energy $\Delta E_{\text{seg}}$ (Section II). The $\Delta F_{\text{seg}}$ distributions of the Mg alloy systems were fitted using a skew-normal function as shown in Figure 2. Compared to the $\Delta E_{\text{seg}}$ distributions, a noticeable shift towards positive values was observed for all solute elements, implying a reduced tendency for segregation as the temperature rises. Despite this shift, all $\mu$ values of the $\Delta F_{\text{seg}}$ distribution remained negative at 600 K, with Nd showing the most negative value at --0.127 eV/atom. For Al, the transition in segregation behavior with temperature is less marked (see Figure 2(e)), as the $\mu$ value of the $\Delta F_{\text{seg}}$ distribution remains consistent with that of the $\Delta E_{\text{seg}}$ distribution, both measured at --0.081 eV/atom. A contraction in the distributions of $\Delta F_{\text{seg}}$ was evident from the decrease in the $\omega$ value, suggesting reduced variability in segregation tendencies at finite temperatures. Notably, the distribution for Ca demonstrated a significant change in skewness at 600 K, with the skewness parameter $\alpha$ shifting from --0.669 at 0 K to a positive value of 0.488. This shift in $\alpha$ indicates a transition in segregation behavior, where favorable segregation sites become less dominant, reflecting a more balanced distribution of segregation tendencies across the GB sites.
@@
Besides quantifying energy spectra, the per-site segregation (free) energy, denoted as $\Delta G_{i}$, can be directly correlated with solute concentration at GBs based on the Langmuir–McLean model, as illustrated in Equation (12) in the context of an infinitesimally dilute solid solution:
@@
where $X_{\text{bulk}}$ is the solute concentration in the bulk region, and $X_{\text{GB},i}$ is the fraction of solute concentration of a GB site $i$. In this work, $\Delta G_{i}$ is referred to as the Gibbs free energy change, which is traditionally used in the Langmuir-McLean model and is approximately equivalent to the Helmholtz free energy change under the simulation conditions, as the pressure-volume work term is negligible in the context of bulk solids. This correlation facilitates the calculation of the weighted average solute concentration at GBs ($X_{\text{GB}}$) as described in Equation (13):
@@
The effect of vibrational entropy on solute segregation at GBs was illustrated by calculating the fractional solute concentration of individual sites according to Equation (12) with and without considering the vibrational term $\Delta F_{\text{seg}}^{\text{vib}}$. The resulting concentrations of various solutes (Nd, Ca, Y, Li, Al, and Zn) at GBs within the simulated polycrystalline Mg at $X_{\text{bulk}}$ of 0.1 at.and 1 at.as a function of temperature are shown in Figure 3. Accounting for the vibrational entropy contribution results in notably lower solute concentrations at GBs than those derived from calculations relying solely on $\Delta E_{\text{seg}}$. The difference in solute concentrations at GBs, attributed to vibrational entropy, becomes increasingly significant with rising temperature and bulk concentration $X_{\text{bulk}}$. Among the solutes studied, Nd stands out for demonstrating a significantly higher $X_{\text{GB}}$ value across the temperature range up to 800 K examined. This selected temperature range slightly exceeds the conventional upper limit of thermomechanical processing temperatures for Mg alloys in experimental settings, which typically extend up to 723 K. This choice ensures that the temperature range of interest is fully covered while allowing for insights into segregation behavior under elevated temperature conditions.
@@ III.3 Machine learning prediction of segregation energy and segregation free energy
Figure 4: Contour plots of ML predicted versus true (a-c) per-site segregation energy $\Delta E_{\text{seg}}$ and (d-f) segregation free energy $\Delta F_{\text{seg}}$ at 600 K of Zn solute segregation at GB sites in the Mg polycrystal. Only the validation data are illustrated here. The CatBoostRegressor model was trained separately using the (a,d) ESD features, (b,e) SOAP features, and (c,f) both ESD and SOAP features. The color density indicates the concentration of data points, with denser regions shown in lighter shades. The diagonal red dashed line represents the ideal scenario where predicted values match the true values perfectly.
@@
Estimating the solute segregation tendency, particularly calculating the vibrational entropy contribution via the HA approach, is computationally expensive for all GB sites in polycrystals. The ML approach is ideal for developing a model to predict the segregation tendency of solutes to GB sites and capture the segregation energy spectrum based on accessible energetic and structural features, which are computationally inexpensive. To predict $\Delta E_{\text{seg}}$, $\Delta F_{\text{seg}}^{\text{vib}}$, and $\Delta F_{\text{seg}}$ at 600 K, we generated 16 ESD that capture the local atomic environment and thermal vibrational properties of GB sites in the polycrystal. The Pearson’s correlation matrices of these features with target segregation energies in different Mg alloy systems are illustrated in Figures S4-9 (Supplementary Material) to identify which features have strong relationships with the target variables and with each other. Among all ESD, hydrostatic stress $\sigma_{h}$, which is calculated based on $\sigma_{xx}$, $\sigma_{yy}$, and $\sigma_{zz}$, shows the highest correlation coefficient above 0.80 with $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ at 600 K for all solute elements, indicating its significant role in determining the GB segregation behavior of solute atoms in the material. The normal stress components $\sigma_{xx}$, $\sigma_{yy}$, and $\sigma_{zz}$ exhibit similar correlation coefficients. Additionally, potential energy, atomic volume, surface area, and coordination number also exhibit high correlation coefficients with the segregation energies.
@@
The contour plots in Figure 4(a) reveal that the predicted $\Delta E_{\text{seg}}$ values align closely with the true values along the diagonal red dashed line, which represents the ideal prediction scenario. The color density represents the concentration of data points, with denser regions displayed in brighter shades. This emphasizes that the highest concentration of data points is near the ideal prediction line, indicating the strong predictive capability of the CatBoostRegressor model with the ESD feature set. The good performance of the CatBoostRegressor model in predicting $\Delta E_{\text{seg}}$ with the ESD features is further demonstrated by the $R^{2}$ scores, with $R^{2}$ between 0.82 and 0.87 for Nd, Y and Al, and between 0.90 and 0.94 for Ca, Li, and Zn, see Figure 4 and Figure S10. In addition, the distributions of the ESD exhibit a high degree of overlap between the predicted and the calculated $\Delta E_{\text{seg}}$, as shown in Figure 5 and Figure S11 for all solute elements. Among all ESD features, $\sigma_{h}$ demonstrates the highest importance, except for Li solute, where the potential energy was identified as the most important feature, see Figure S12. Potential energy was ranked as the second most important feature for Al and Zn solute elements and the third for Ca. The structural features obtained using Voronoi tessellation, including atomic volume, surface area, and cavity radius, were ranked highly in importance, e.g., for Nd and Zn solute elements, atomic volume is the second most important feature, while for Y, cavity radius was ranked second. The vibrational properties calculated from the series of MD simulations, including vibrational MSD $MSD_{\text{vib}}$ and flexibility volume $V_{\text{flex}}$, were demonstrated as important features in predicting $\Delta E_{\text{seg}}$: for Nd, Ca, Y, Al, and Zn, these vibrational features are among the top five most important features.
@@
Figure 5: Distributions of ML predicted and true (a-c) per-site segregation energy $\Delta E_{\text{seg}}$ and (d-f) segregation free energy $\Delta F_{\text{seg}}$ at 600 K of Zn solute segregation at GB sites in the Mg polycrystal. Only the validation data are illustrated here. The CatBoostRegressor model was trained separately using the (a,d) ESD features, (b,e) SOAP features, and (c,f) both ESD and SOAP features. Distributions of predicted versus true energies with a bin size of 0.01 eV/atom.
@@
In addition to the ESD features, we generated SOAP features for automatically describing the fingerprint of the local atomic environment at GB sites in the polycrystal. These SOAP descriptors were fed to the CatBoostRegressor model to predict $\Delta E_{\text{seg}}$, $\Delta F_{\text{seg}}^{\text{vib}}$, and $\Delta F_{\text{seg}}$ at 600 K (Figure 4-5 and Figures S19-24). The SOAP-featured model shows better performance in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ at 600 K than the ESD-featured model for all solute elements, except for Li, as demonstrated by higher $R^{2}$ scores and lower MAE (see Figure 6). However, the performance in predicting $\Delta F_{\text{seg}}^{\text{vib}}$ remains poor (Figures S21,22). By merging ESD and SOAP descriptors, the ML model shows further improvement in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$, as evidenced by the tighter clustering around the diagonal (Figure 4(c,f)) and with the $R^{2}$ score of all solute element in predicting $\Delta E_{\text{seg}}$ above 0.85 and $\Delta F_{\text{seg}}$ above 0.80 (Figures S25-30). However, the improvement by utilizing combined ESD and SOAP features in predicting $\Delta F_{\text{seg}}^{\text{vib}}$ is limited. The model’s performance in predicting $\Delta F_{\text{seg}}$, which reflects a rule of mixture, combining the high accuracy of $\Delta E_{\text{seg}}$ predictions with the relatively lower accuracy of $\Delta F_{\text{seg}}^{\text{vib}}$ predictions. This behavior arises because $\Delta F_{\text{seg}}$ is a linear combination of $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}^{\text{vib}}$ as illustrated in Equation (4).
The potential redundancy and overfitting in the training features were evaluated. For instance, some of the ESDs may contain overlapping information, such as atomic volume being related to both hydrostatic stress and flexibility volume. However, the performance metrics of the model across training, testing, and unseen validation datasets indicate only slight differences, which suggests that overfitting is not occurring, as demonstrated in Table S2. Additionally, a learning curve analysis by plotting the training and validation loss against the size of the training dataset shows an initial low training error, which gradually increases and then stabilizes as the training size grows, see Figure S31. This behavior indicates that the model tends to memorize patterns with smaller datasets, resulting in low training error. However, as more data is introduced, the model shifts from memorization to generalization, leading to a rise in training error. Eventually, the training error plateaus when the model reaches its learning capacity, meaning additional data does not significantly impact performance. The comparable validation errors further confirm that the model is well-suited for this study and does not suffer from overfitting. Furthermore, the feature overlap does not appear to significantly impact predictive accuracy as the effect of dimensionality reduction by removing highly correlated ESD features (correlation threshold = 0.9) did not result in any significant improvement in model performance, with both MAE and $R^{2}$ values remaining nearly identical, see Table S3.
For the SOAP features, dimensionality reduction using Principal Component Analysis (PCA) or feature selection techniques by selecting the top-20 important SOAP features based on feature importance rankings were applied. As shown in Table S4, neither using 20 PCA components nor restricting the model to the top-20 important features resulted in improved performance compared to the full set of SOAP features. The MAE and $R^{2}$ values for these reduced models indicate a slight decrease in prediction accuracy. While dimensionality reduction did not enhance accuracy, it remains a valuable tool for improving the model’s interpretability, particularly in scenarios where computational resources or explainability are key considerations.
@@
The residual distributions of predicted and true $\Delta E_{\text{seg}}$, $\Delta F_{\text{seg}}^{\text{vib}}$, and $\Delta F_{\text{seg}}$ via ESD, SOAP, and combined ESD and SOAP features are shown in Figures S32-34. A left-skewed residual distribution in $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ was observed for all solute elements and feature sets, indicating the models consistently over or under-predict the solute segregation tendency at favorable segregation sites. For future work, uncertainty determination will be crucial to better understand and address these biases and refine the models accordingly.